GreenHab Report – March 17th

GreenHab Officer: Diane Rothberg

Environmental control: Heating and cooling w/ambient air (8 hrs)

Shade cloth (40%)

Average temperatures:

Low: 15 ° C

High: 32 ° C

Hours of supplemental light: 0 hours

Daily water usage for crops: 13 gal

Water in Blue Tank—gallons 181 gal

Time(s) of watering for crops: 8:30 16:45

Changes to crops:

Narrative: Discarded 4 pots of dead peas, but kept the 2 best just in case they revived

Harvest: 19 gm basil

Support/supplies needed: Getting low on fly paper sheets

Operations Report – March 17th

SOL: 1

Name of person filing report: Tom Baldwin

Non-nominal systems: Rover charging system.

Notes on non-nominal systems: Circuit to charging system interrupted, reset by David Mateus

Generator (hours run): 11

Solar— 81 % (Before generator is run at night)

Diesel Reading – 65 %

Propane Reading – 68 %

Ethanol Free Gasoline – gallons.

Water (auxiliary tank) – 0 gallons

Water (static tank) – 350 gallons

Auxiliary to Static tank transfer– no

Gallons transferred:

Water in GreenHab – 181 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 12 units

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: yes

Hours: 83.6

Beginning charge: 48

Ending charge: 46

Currently charging: yes

Opportunity rover used:


Beginning charge: 58

Ending charge: 56

Currently charging:yes

Curiosity rover used: yes

Hours: 82.9

Beginning charge: 100

Ending charge: 98

Currently charging: yes

Notes on rovers: Demo/instruction by Dave Mateus

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) no

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? no

CrewCar used and why, where? no

General notes and comments: nothing to report

Summary of internet: nothing to report

Summary of suits and radios: nothing to report

Summary of Hab operations: Clean up, cleaning, & report writing

Summary of GreenHab operations: Monitoring $ watering of plants

Summary of ScienceDome operations: cleaned & organized

Summary of RAMM operations: nothing to report

Summary of any observatory issues: Power bar switch was intermittent, now normal.

Summary of health and safety issues: Check of emergency equipment. 1 smoke detector low battery

Questions, concerns and requests to Mission Support: none

Astronomy Report – March 17th

Astronomy Report

Name: Tom Baldwin

Crew: 207

Date: 17-MAR-2019


Robotic Telescope Requested MDRS-WF

Objects Viewed: Crab Nebula

Problems Encountered:

MUSK OBSERVATORY Dry run on set-up only

Solar Features Observed: none

Problems Encountered: Power bar was intermittent. Seems normal now. Mentioned to David Mateus.

Sol Summary – March 17th

Summary Title: Starting our Sim

Author: Julielynn Wong

Mission Status: All nominal

Sol Activity Summary: Crew 206 departed this morning. Crew 207 Greenhab Officer prepared breakfast. Crew 207 Commander submitted yesterday’s food inventory. Crew 207 HSO conducted a safety check. Crew 207 Scientist drafted the cooking and cleaning schedule. Commander cleaned and organized both levels of the Hab and the Science Dome. Greenhab Officer and Crew 207 Engineer/Astronomer began planning tomorrow’s EVA. Crew Scientist prepared lunch with fresh basil seasoning from the Greenhab. Engineer/Astronomer began operations in the solar observatory. Crew 207 completed paperwork and training with the MDRS intern. All 5 crew members successfully test drove the rovers under the supervision of the MDRS intern. Crew 207 completed their EVA spacesuit training.

Crew 207 identified the following useful items that could be 3D printed out of biodegradable, washable plastic at MDRS:

Coat hanger

Bag clip

Soap holder

S hook

Stackable storage bin

Pen storage container


Labware drying rack

Soil sieves

Water dipstick

Selfie stick


Colored straws to help crew members identify their individual cups

Scratch stick for the inside of the EVA spacesuit helmets

Look ahead plan: We are planning an EVA to scout out a location for drone flight testing. We are preparing files to be 3D printed at MDRS to save time and money.

Anomalies in work: None

Weather: Sunny all day

Crew physical status: Fine


Reports to be filed: Operations, GreenHab, HSO, Scientist, Journalist, EVA Request

Support Requested: None

Journalist Report – March 17th

Author: Diane Rothberg

We spent this beautiful, sunny day out of sim, as we’re waiting for so much to arrive…3D printers, drones, a bike, and two additional crew members.

Still getting the hang of water conservation, which is tricky when you have to wash a sticky oatmeal pot, as well as last night’s pots holding clingy, aging leftover soup and rice.

Today’s lunch of spaghetti bolognaise prepared by Tiffany and Kevin was a huge hit, especially after we supplemented it with fresh basil from the Greenhab.

Our commander, Julielynn, did some much-needed cleaning and organizing around the hab and the science dome, and came up with creative ideas as to what items we should 3D print to make life easier here on Mars.

Kevin, our health and safety officer checked all CO2 detectors and fire alarms.

We received a safety briefing from David, and we all got rover check out and a chance to drive them before we have to do it in a bulky space suit.

Late in the afternoon, we were checked out on the space suits, and now, our Mars simulation truly begins!

Copyright © The Mars Society. All rights reserved. | Main Site